OpenStudy (anonymous):

2y 8 -- = -- 9 4y PLEASE I NEED HELP WORKING ON MY FINALS!! :(

OpenStudy (anonymous):

cross multiply and u will get a quadratic equation

OpenStudy (anonymous):

can you help me through it? Im so bad with math

OpenStudy (anonymous):

yes..i 'll try to help u..in these type of problems...multiply the numerator of the first fraction and denominator of the second fraction and multiply the DENOMINATOR of the first fraction and NUMERATOR of the second fraction AND PUT AN EQUAL SIGN BETWEEN THEM

OpenStudy (anonymous):

is it 6 and -6?

OpenStudy (anonymous):

now find the value of y

OpenStudy (anonymous):

sry will get8y^2=72

OpenStudy (anonymous):

how do i get y's value? do i multiply

OpenStudy (anonymous):

8*y^2=72........divide both the sides of the equation by 8........u will get y^2=9..now find the square root of both the sides....u will get y=3 and -3

OpenStudy (anonymous):

ok, im sorry i dont get things very well, may i ask you one more question?

OpenStudy (anonymous):

ok..ask

OpenStudy (anonymous):

XYZ is a triangle with side lengths in the ratio 2 : 3 : 4. The perimeter of XYZ is 27 yards. Find the length of the longest side.

OpenStudy (anonymous):

ill put the picture in a sec

OpenStudy (anonymous):



OpenStudy (anonymous):

since the sides are given in ration.. u can form a linear equation to find the longest side...let me give u a step by step process

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

we know that perimeter means the sum of all the sides......so here sum of all the sides is 4x+2x+3x..again it is given that the perimeter of the triangle is 27...so we can write....4x+2x+3x=27... got it?

OpenStudy (anonymous):

ok im following along as best i can

OpenStudy (anonymous):

now can u find the value of x from the above equation?

OpenStudy (anonymous):

i dont know if i remember how to solve that? can you show me the steps and ill learn as we go?

OpenStudy (anonymous):

sry i need to go now.....plz help him someone....@thomaster @Abhisar @tkhunny

OpenStudy (anonymous):

oh ok, well thank you for helping me! and im a girl :) lol

OpenStudy (anonymous):

oh..sry..by bye...i hope someone will come and help u

OpenStudy (tkhunny):

\(\dfrac{2y}{9} = \dfrac{8}{4y}\) \(\dfrac{y}{9} = \dfrac{4}{4y}\) \(\dfrac{y}{9} = \dfrac{1}{y}\) In this last form, the solution is obviously 3. It is unlikely that in such a proportion a negative number would be an appropriate result - particularly if you are going to draw triangles to explain it. \(y = \dfrac{9}{y}\) \(y^{2} = 9 \rightarrow y = 3\)

OpenStudy (anonymous):

Oh thank you so much!! Could you help me with a couple more problems?

OpenStudy (tkhunny):

Post a new thread.

OpenStudy (anonymous):

A circle has a circumference of 14 ft. Its area is halved. Describe the effect on the circumference of the circle. The circumference is divided by √2. The circumference is divided by 2. The circumference doubles. The circumference is divided by √3.